OpenStudy (annie96):

What is the value of f(-4)? f:x-->3-x A.-1 B.7 C.-4 D.3

OpenStudy (anonymous):

-1 lol

jimthompson5910 (jim_thompson5910):

f(x) = 3-x f(-4) = 3 - (-4) ... replace every x with -4. Now simplify

OpenStudy (annie96):

@jim_thompson5910 is it 7?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (annie96):

thank you

jimthompson5910 (jim_thompson5910):

since 3 - (-4) = 3+4